OpenStudy (anonymous):

When using the substitution method, what expression for x in the first equation would be substituted in the second equation? x + 3y = -7 2x + 4y = -4

OpenStudy (anonymous):

Okay, so basically, in the first equation, get x alone so that it's in the form: x = stuff So what's the first step in doing that?

OpenStudy (anonymous):

would i subtract it from both sides?

OpenStudy (anonymous):

Okay, so what did you get?

OpenStudy (anonymous):

Am I getting rid of the x or the y? If i did x it would it be 3y = -7 - x ?? Im not sure how to do this..

OpenStudy (anonymous):

or 3y = -6?

OpenStudy (anonymous):

You want it to be x = stuff

OpenStudy (anonymous):

x + 3y = -7 So we're just going to subtract the 3y from the left side so that x is alone. x+3y-3y = -7-3y x=-7-3y So in the second equation, everywhere you see and x, put -7-3y instead.

OpenStudy (anonymous):

after you wrote that I realized i was supposed to subtract the 3y, I just get confused on that part.. Thank you!

OpenStudy (anonymous):

2nd equation: 2x + 4y = -4 So if you were solving it, you would find the x and get rid of it, in its place put -7-3y 2(-7-3y) + 4y = -4 Distribute the two: -14 - 6y + 4y = -4 Add -6y to 4y -14 - 2y = -4 Work on getting y alone, so we'll add 14 to that side to get rid of the -14 (and add it to the right side of the equation so it's still equal.) -14+14 - 2y = -4+14 -2y=10 divide off the -2 (-2y)/-2=10/-2 y = -5 So now we know y is equal to -5. So in either of the original two equations (whichever one you dislike the least), substitute -5 for y and solve it so you can find out what x really is. I'll use the first equation for this: x + 3y = -7 x + 3(-5) = -7 x + -15 = -7 x + -15 + 15 = -7 + 15 x= 8 So in your system of equations, y = -5 and x = 8